# Monocular Facial Appearance Capture in the Wild

### v1

@@ 4.2 Comparisons
We also perform qualitative comparisons to NextFace and SunStage in Fig. 4. NextFace relies on a statistical prior, leading to inaccurate shapes and blurry appearance for different subjects (Fig. 4, rows 1 and 4). SunStage assumes a single point light in the scene, and the shadows and specular components come only from this point light. This leads to poor diffuse and specular separation and incorrect shadows in generic environments like the examples shown here (Fig. 4, rows 2 and 5). In the first subject, we see orange artifacts on the forehead of the diffuse albedo in the SunStage result (row 2). In the second subject, the specular component is completely missing (row 5). The shape and textures from SunStage also have lower resolution than ours. The final row of Fig. 4 illustrates the appearance details for zoomed-in regions shown by the red and blue squares, indicating that our method produces the most accurate details.
@@
4.3 Ablations
Raytracing vs pre-filtered environment map for diffuse.
@@ 4.3 Ablations
We first compare ray tracing versus pre-filtered environment map in terms of joint reconstruction of diffuse albedo, geometry and illumination. The specular component is discussed separately in the next paragraph as it contributes mostly to the micro- instead of the macro-geometry. We created a synthetic dataset with a monocular head rotation sequence similar to the real data. The ground truth mesh and albedo are from a studio appearance capture method [52] (Fig. 5 column 1). We use a natural outdoor environment map on a cloudy day as lighting. Please refer to the supplementary document for more details on the synthetic dataset. We show reconstructed diffuse albedo and geometry of ray tracing and pre-filtered environment map in columns 2 and 3 of Fig. 5, and the error maps compared with the ground truth are shown in the last two columns. We can see that ray tracing produces a more accurate reconstruction of the diffuse albedo while the result from the pre-filtered environment map contains a lot of baked-in shading. In terms of shape reconstruction, the two perform similarly. However, the pre-filtered environment map does not handle shadows well, leading to slightly worse shape recovery in regions where self-occlusion plays an important role, e.g., the lips.
Raytracing vs pre-filtered environment map for specular.
@@
The original render from the pre-filtered environment map has some unrealistic highlights at the grazing angle on the side of the nose when inspecting the specular component (Fig. 6, row 1 column 2). This is because the shading model does not account for light visibility and shows the contribution of the light from behind the head. Our proposed view-dependent visibility attenuation (Fig. 6, row 1 column 1) effectively masks out these regions and the final render matches the ray traced version closely. However, in our inverse rendering setting when we need to solve for shape, textures, lighting at the same time, ray tracing produces flickery images (Fig. 6 row 2) where part of the skin abruptly changes from very bright to very dark while the head is rotating. We also notice high frequency artifacts, e.g. sharp boundary between bright and dark pixels, in the ray traced render, as denoted by the red arrows. In contrast, our visibility modulated pre-filtered environment maps used for the specular component give visually smoother specular renders (Fig. 6, row 3). Note that the ray traced and pre-filtered environment map results are rendered from different sets of optimized lighting and textures in this example so the specular highlights do not match exactly.
@@
Figure 5: Comparing ray tracing and pre-filtered environment map on a synthetic dataset rendered with a Lambertian diffuse material. The mesh errors are displayed with a scale of 0mm Refer to caption 5mm and the albedo error with a scale of -0.1 Refer to caption 0.1.
Figure 6: Comparing the specular renders from ray tracing and a pre-filtered environment map. Our proposed view-dependent specular visibility term correctly masks out the unrealistic highlights on the side of the nose, resembling the ray traced results, while produces smoother renders.
@@
We apply a weak regularization to encourage the diffuse render to be small which prevents too much specular from being baked into the diffuse component, as we show in Fig. 7. This gives us better separation of the diffuse and specular signals.
@@
Optimizing vertex positions vs blendweights.
Last, we choose to optimize vertex positions directly instead of parameterizing the shape using blendweights of a 3DMM as in NextFace [16]. We show in Fig. 8 that our model achieves the least reconstruction error in the central part of the face compared to the initial mesh or the one optimized through a 3DMM in the same inverse rendering setting. Even though the silhouette of the side face is never shown in the training data, our method recovers the correct shape of the nose from shading. The ground truth reference mesh is a 3D facial scan from a multi-view face scanner [4].
Figure 8: Our geometry optimization pipeline recovers the correct nose shape from shading, achieving the better reconstruction quality compared to 3DMM fit. The mesh errors are displayed with a scale of 0mm Refer to caption 5mm
@@ 6 Conclusion
We propose a new lightweight facial appearance capture method, surpassing the quality of existing lightweight approaches. Our method works truly in the wild, making no assumption on the environment illumination. Our novel shading model that combines ray tracing and a visibility modulated environment map leads to faithful recovery of the shape and appearance properties from a monocular video containing a simple head rotation.
@@
In this supplementary document we begin by discussing the details of our capture protocol and initial tracking steps in Section 7. In Section 8 we highlight the particulars of our implementation of the baseline methods we use for comparisons, and elaborate on the synthetic dataset we used for evaluation. Finally, we add additional results and failure cases in Section 9.
@@ 7.2 Initial Tracking
In the initial tracking stage, we estimate an initial mesh and per frame head poses. The initial mesh is parameterized using blendweights of a PCA face basis computed from the dataset of Chandran et al. [11], which includes 50 eigen faces for identity and 25 for expression. Our tracking algorithm uses a combination of a landmark [12] loss and a photometric loss, similar to Qian [50]. The only difference is that we solve for a global expression code since we assume the expression does not change in the same sequence. We apply a weight of 100 on the landmark loss and a weight of 30 on the photometric loss for all our datasets.
@@ 8 Experiment Details
8.1 Baseline Implementation
Next we describe the steps performed to run the baseline comparisons. We use the default parameters of FLARE [5]. We noticed that the FLARE geometry is very bumpy, hence we tried setting a larger weight on the Laplacian mesh regularizer. However, this resulted in flatter face geometry and did not improve the results.
@@ 9 Additional Results
Visualization of specular visibility.
First we provide additional visualizations of the view-dependent specular visibility under different roughness in Fig. 11. This highlights the areas that are impacted by the visibility computation. When the roughness is small (mirror material), this visibility term is close to a binary mask and when the roughness is large (diffuse material), it gets closer to a view-independent ambient occlusion term. Note that the approximation error of Eq. 7 is small for a smaller roughness value and big for a larger roughness value. Applying the same approximation for the diffuse component would lead to a very large error.
Figure 11: Visualization of the estimated view-dependent specular visibility term with difference roughness values.
9.1 Failure Cases

### v2

@@ 4.2 Comparisons
We also perform qualitative comparisons to NextFace and SunStage in Fig. 4. NextFace relies on a statistical prior, leading to inaccurate shapes and blurry appearance for different subjects (Fig. 4, rows 1 and 4). It produces very similar shapes for the two subjects shown in Fig. 4, while our geometry preserves the identity and likeness of the subjects. SunStage assumes a single point light in the scene, and the shadows and specular components come only from this point light. This leads to poor diffuse and specular separation and incorrect shadows in generic environments like the examples shown here (Fig. 4, rows 2 and 5). In the first subject, we see orange artifacts on the forehead of the diffuse albedo in the SunStage result (row 2). In the second subject, the specular component is completely missing (row 5). The shape and textures from SunStage also have lower resolution than ours. The final row of Fig. 4 illustrates the appearance details for zoomed-in regions shown by the red and blue squares, indicating that our method produces the most accurate details.
@@
4.3 Evaluations
Effects of light visibility.
We first evaluate the effects of accounting for self-occlusion by comparing the relit renders of our method with the original split-sum approximation (denoted as Ours (w/o vis)) in Fig. 5. We can see that when self-occlusion is not accounted for, shadows under the capture lighting got baked into the albedo (column 1 with zoomed-in patches focusing around the nose), leading to wrong shadows when relit (row 1, column 2, 3). The specular component also exhibits unrealistic sharp highlight (row 1, column 2, 4). Moreover, when the major light source is behind the subject, the baseline renders show artifacts in the form of strong specular highlights on the side of the nose and under the chin (row 3, column 2, 4) since self-occlusion is not properly handled. In contrast, our proposed model correctly removes baked-in shading and produces more realistic relit renders (row 2, 4). We also quantitatively evaluated the performance gain of accounting for self-occlusion in terms of reconstruction errors in Table 1.
Figure 5: Relighting evaluation of our method with and without accounting for self-occlusion. We also visualize the corresponding diffuse albedo, diffuse render and specular render.
Evaluation on synthetic data.
@@ 4.3 Evaluations
To better evaluate the quality of the reconstructed mesh and textures against ground truth, we created a synthetic dataset with a monocular head rotation sequence similar to the real data. The ground truth mesh and albedo are from a studio appearance capture method [55] (Fig. 6 column 1). We use a natural outdoor environment map on a cloudy day as lighting. Please refer to the supplementary document for more details on the synthetic dataset. We show reconstructed diffuse albedo and geometry of our method with and without light visbility accounted for in columns 2 and 3 of Fig. 6, and the error maps compared with the ground truth are shown in the last two columns. We can see that our method produces a more accurate reconstruction of the diffuse albedo while the baseline result contains a lot of baked-in shading. In terms of shape reconstruction, the two perform similarly. However, the baseline method does not handle shadows well, leading to slightly worse shape recovery in regions where self-occlusion plays an important role, e.g., the lips.
Ours vs raytracing for specular.
@@
While a ray-tracer handles visibility inherently, we find that our modified split-sum approximation behaves more stably in our inverse rendering setting when we need to solve for shape, textures, lighting at the same time from merely monocular input. Ray tracing produces flickery specular images (Fig. 7 row 2) where part of the skin abruptly changes from very bright to very dark while the head is rotating. We also notice high frequency artifacts, e.g. sharp boundary between bright and dark pixels, in the ray traced render, as denoted by the red arrows. In contrast, our model gives visually smoother specular renders (Fig. 7, row 3). Please refer to the supplemental videos for an animated visualization.
@@
Figure 6: Comparing our method with and without accounting for light visibility on a synthetic dataset. The mesh errors are displayed with a scale of 0mm Refer to caption 5mm and the albedo error with a scale of -0.1 Refer to caption 0.1.
Figure 7: Comparing the specular renders from our method and a ray-tracer across frames in the same head rotation sequence.
Optimizing vertex positions vs blendweights.
We choose to optimize vertex positions directly instead of parameterizing the shape using blendweights of a 3DMM as in NextFace [15]. We show in Fig. 8 that our model achieves the least reconstruction error in the central part of the face compared to the initial mesh or the one optimized through a 3DMM in the same inverse rendering setting. Even though the silhouette of the side face is never shown in the training data, our method recovers the correct shape of the nose from shading. The ground truth reference mesh is a 3D facial scan from a multi-view face scanner [3].
Figure 8: Our geometry optimization pipeline recovers the correct nose shape from shading, achieving the better reconstruction quality compared to 3DMM fit. The mesh errors are displayed with a scale of 0mm Refer to caption 5mm.
@@
Last, we apply a weak regularization to encourage the diffuse render to be small which prevents too much specular from being baked into the diffuse component, as we show in Fig. 9. This gives us better separation of the diffuse and specular signals.
@@ 6 Conclusion
We propose a new lightweight facial appearance capture method, surpassing the quality of existing lightweight approaches. It works truly in the wild, making no assumption on the environment illumination. Our novel shading model explicitly accounts for self-occlusion, leading to faithful recovery of the shape and appearance properties from a monocular video containing a simple head rotation.
@@
In this supplementary document we begin by discussing the details of our capture protocol and initial tracking steps in Section 7. We provide more implementation details in Section 8. In Section 9 we highlight the particulars of our implementation of the related methods we use for comparisons, and elaborate on the synthetic dataset we used for evaluation. Finally, we add additional results and failure cases in Section 10.
@@ 7.2 Initial Tracking
In the initial tracking stage, we estimate an initial mesh and per frame head poses. The initial mesh is parameterized using blendweights of a PCA face basis computed from the dataset of Chandran et al. [10], which includes 50 eigen faces for identity and 25 for expression. Our tracking algorithm uses the combination of a landmark [11] loss and a photometric loss, similar to Qian [53]. The only difference is that we solve for a global expression code since we assume the expression does not change in the same sequence. We apply a weight of 100 on the landmark loss and a weight of 30 on the photometric loss for all our datasets.
@@
8 Implementation Details
We capture 500 to 800 frames for each subject. We then uniformly sample around 250 frames to use in the inverse rendering. We did not observe a performance gain or drop when using all of the frames. The images are cropped to 1K resolution. We also solve the texture maps at 1K resolution. The environment map is a cubemap with resolution $6\times 256\times 256$ at the largest mip-level, and resolution $6\times 8\times 8$ at the smallest mip-level. For each pixel, we draw $256$ light samples and $256$ BRDF (cosine) samples for the diffuse render, and $64$ samples to estimate the view-dependent specular visibility. We employ the Adam optimizer [19] with a learning of $0.1$ for vertex positions and the environment map, and $0.001$ for the textures. We use the differentiable rasterizer from Laine et al. [37] to obtain the primary visibility and the OptiX [52] engine for ray tracing. Each subject is trained for $6000$ iterations which takes around 2 hours on a Nvidia RTX 3090 GPU. The weights for Eq. 10 are defined as
$0.1,\displaystyle\lambda_{\text{mask}}=1,\lambda_{\text{Lap}}=0,\lambda_{\text{light}}=$1, (11)
[…]
@@ 9 Experiment Details
9.1 Implementation of the Related Methods
Next we describe the steps performed to run the comparisons. We use the default parameters of FLARE [4]. We noticed that the FLARE geometry is very bumpy, hence we tried setting a larger weight on the Laplacian mesh regularizer. However, this resulted in flatter face geometry and did not improve the results.
@@ 10 Additional Results
Visualization of the Visibility.
First we provide additional visualizations of the view-dependent visibility under different roughness in Fig. 12. This highlights the areas that are impacted by the visibility computation. When the roughness is small (mirror material), this visibility term is close to a binary mask and when the roughness is large (diffuse material), it gets closer to a view-independent ambient occlusion term. Note that the approximation error of  Eq. 6 is small for a smaller roughness value and big for a larger roughness value. Applying the same approximation for the diffuse component would lead to a large error. We also show an example in Fig. 13 of how our visibility-modulated split-sum approximation can be used in other rendering tasks as a practical way to add self-shadowing to glossy objects.
Figure 12: Visualization of the estimated view-dependent specular visibility term with different roughness values.
Figure 13: Rendering of a glossy Happy Buddha.
Failure Cases.
@@
Effects of $\lambda_{geo}$.
We use $\lambda_{geo}=19$ based on [50], but this smoothing term can be tuned for different subjects to get better geometry. An ablation for different values of $\lambda_{geo}$ is shown in Fig. 16. Note that a large $\lambda_{geo}$ leads to loss of details, but a small $\lambda_{geo}$ can lead to self-intersections.
Figure 16: Effects of $\lambda_{geo}$. The shaded meshes are shown on the top row with the corresponding normals visualized on the bottom.
Novel Views in the Capture Environment.
We show novel view renders in the capture environment in Fig. 17. We primarily focused on the facial regions for this project. Artifacts around the boundaries can be improved with some engineering efforts, such as better masking and better initial alignment. Reconstruction of the hair and shoulders is also an interesting area for future work, which can improve the overall quality.
Figure 17: Novel views rendered in the capture environment (top row) along with the corresponding meshes (bottom row).
Comparison with CoRA.
CoRA [28] requires a more constrained capture setup with a co-located light and camera in a dark room. We ran our method and CoRA on a subject using the CoRA capture protocol, see Fig. 18, demonstrating that our method achieves equally good reconstruction given the same data. Additionally, our method also works in more generic environments.
Figure 18: Comparison with CoRA on their data protocol.